• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Spiegel

    Log entry time 11:43:24 on August 17, 2006

    Entry number 180524

    keyword=Bogie Repair

    Mark and I performed a follow-up diagnosis of the bogie failure from yesterday. Mark re-connected the drive coupling while I replaced the motor encoder. Tested drive system in CW and CCW directions for 2 degrees of traval. Everything worked smoothly. Re-postioned LEFT spectometer to 92.831 degrees.

    A copy of this log entry has been emailed to: kees@jlab.org,stevensm@jlab.org